OpenStudy (anonymous):

can you please help me with this question please Find the probability of the following five-card poker hands from a 52-card deck. In poker, aces are either high or low. Two pair (2 cards of one value, 2 of another value) A. 20592/433115 B. 22464/433115 C.3432/433115 D. 5148/433115

OpenStudy (anonymous):

we can work through this i am sure, but it is probably better to just google this all stuff about poker on the web

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Poker_probability

OpenStudy (anonymous):

I did am like 3 hours trying to figure it out

OpenStudy (anonymous):

ok maybe it would be easier if we figured out what the probability of getting a pair of 2 and a pair of 3 it would be \[\frac{\binom{4}{2}\binom{4}{2}\binom{44}{1}}{\binom{52}{5}}\] i think

OpenStudy (anonymous):

reasoning 2 out of the 4 twos 2 out of the 4 threes and one out of the remaining cards that are neither 2 nor 3

OpenStudy (anonymous):

but how do I get the answer?

OpenStudy (anonymous):

is it more or less clear how i got the answer above?

OpenStudy (anonymous):

we are not done, but almost

OpenStudy (anonymous):

i computed it for one particular pair, a pair of twos and a pair of threes there are 13 face values in the deck, and you are picking two of them so we have to multiply by \(\binom{13}{2}\) to get the final answer it should be \[\frac{\binom{13}{2}\binom{4}{2}\binom{4}{2}\binom{44}{1}}{\binom{52}{5}}\]

OpenStudy (anonymous):

are you asking how to compute that number?

OpenStudy (anonymous):

2*13= 26

OpenStudy (anonymous):

am confuse

OpenStudy (anonymous):

\(\binom{13}{2}=\frac{13\times 12}{2}=78\)

OpenStudy (anonymous):

\(\binom{44}{1}=44\) and \[\binom{4}{2}=\frac{4\times 3}{2}=6\]

OpenStudy (anonymous):

so your numerator is \[78\times 6\times 6\times 44=123552\]

OpenStudy (anonymous):

denominator is \[\binom{52}{5}=2598960\]

OpenStudy (anonymous):

final answer is \[\frac{123552}{259862}\] or whatever you get when you reduce that fraction

OpenStudy (anonymous):

it's not close to any of the answers giving in A B C or D

OpenStudy (anonymous):

i am going to stick with that answer especially since that is what wiki says it should be

OpenStudy (anonymous):

wiki answer says frequency is \(123552\) and the denominator has to be \(\binom{52}{5}=2598960\)

OpenStudy (anonymous):

Yea but am suppose to choose one of the above

OpenStudy (anonymous):

here is the reduces fraction you see the percent is what wiki says it should be http://www.wolframalpha.com/input/?i=123552%2F2598960

OpenStudy (anonymous):

do I have to sing up to see how I can work the problem?

OpenStudy (anonymous):

i'll just guess

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

Am smarter than you I chose A and I got it correct yea meeee